A mountain resort that featured skiing in the winter was partially staffed by college students in the summer. One summer there were 3 times as many students employed as there were year-round employees. When September came, 40 of the students were back to school and 30 nonstudents were hired for winter. If there were the twice as many nonstudents as students, how many people staffed the resort in the winter?
Answer by Abbey(339)

let x=number of non-students in summer
let 3x-number of students in summer (3 times as many)
3x-40 = number of students in winter (40 went back to school)
x+30 = nonstudents in winter (30 more were hired)
if there were twice as many nonstudents then:
x+30=2(3x-40)
x+30=6x-80
x+110=6x
110=5x
22=x
now verify:
3x = 66 summer students
if 40 left = 26 winter students
x+30=52 winter non students, and this is twice as many as 26, so the answer is correct.
In the winter 26+52=78 people staffed the resort.